Two angles of a triangle have the same measure and the third one is 3 degrees greater than the measure of each of the other two. Find the measure of the LARGEST angle in the triangle.
Answer: The LARGEST angle has a measure of _________ degrees.

Answer by algebrahouse.com(1659)

x = measure of one angle
x = measure of other angle {two angles are equal}
x + 3 = measure of the third angle {third is 3° greater than each of the other two}

x + x + x + 3 = 180° {angles of a triangle add up to 180}
3x + 3 = 180 {combined like terms}
3x = 177 {subtracted 3 from each side}
x = 59 {divided each side by 3}
x + 3 = 62 {substituted 59, in for x, into x + 3}

62° is the measure of the larger angle

For more help from me, visit: www.algebrahouse.com